OpenStudy (anonymous):

Which two lines are parallel? I. 5y = –3x – 5 II. 5y = –5 – 3x III. 3y – 2x = –1

OpenStudy (callisto):

Rearrange the equation into a slope-intercept form: y=mx+c, where m=slope I. 5y = –3x – 5 => y = -3/5 y -1 II. 5y = –5 – 3x => y = -3/5 y -1 III. 3y – 2x = –1 => 3y = 2x-1 => y= 2/3 x -1/3 So can you figure the answer out?

OpenStudy (anonymous):

a and b?

OpenStudy (callisto):

yea :)

OpenStudy (anonymous):

thank you(:

OpenStudy (callisto):

welcome, do you understand how to do this type of questions now?

OpenStudy (anonymous):

yes i do, thank youuuu!

OpenStudy (callisto):

You're welcome :) Glad to hear :)

OpenStudy (anonymous):

Whenever you need to evaluate if things are perpendicular or parallel, get them into the y=mx+b format and look at the values of m. If they are the same, they are parallel. If they are negative reciporacals, they are perpendicular :)